Solve the given homogeneous differential equation
X2dy+y(x+y)dx=0
1
Expert's answer
2020-12-16T14:52:28-0500

"x^2dy+y(x+y)dx=0"
1) Expand:x² dy+yx dx +y² dx = 0or"y'+ \dfrac{y}{x}+ \dfrac{y^2}{x^2}=0" .....(*)   first-order nonlinear ODE.


2) We need to transform it into linear ODEObserve the exponent of y, it's n=2. So we'll use substitution "v=y^{1-n}=y^{1-2}=y^{-1}"
3) So we have"v= \dfrac{1}{y}" "v'= (\\dfrac{1}{y})\t'" "v'=-( \\dfrac{1}{y^2}\t)y'=-v^2y'" In (*) substitute "y'=-( \\dfrac{1}{v^2}\t)v'" and "y= \dfrac{1}{v}" "\\dfrac{-v}{v^2}\t+ \\dfrac{1}{xv}+ \\dfrac{1}{x^2v^2}=0" .... (**)
this is a first-order linear ODE.4) New substitutionv = x uwhere u=u(x) and that's all we need to know about "u" for nowv' = xu' + uback to (**)"- \\dfrac{(xu'+u)}{x^2u^2}\t+ \\dfrac{1}{x^2u}\t+ \\dfrac{1}{(x^4u^2)}\t=0" or expanded"- \\dfrac{u'}{xu^2}\t- \\dfrac{1}{x^2u}\t+\\dfrac{1}{(x^2u)}+ \\dfrac{1}{(x^4u^2)}\t=0" 5) Solve for u' 
two middle terms are canceled:"-\dfrac{u'}{(xu^2)}+\dfrac{1}{(x^4u^2)}=0"
multiply by xu²"-u'+ \\dfrac{1}{x^3}\t=0" "u'= \dfrac{1}{x^3}"
6) Integrate with respect to x"u=\\int( \\dfrac{1}{x^3}\t)dx" "u=- \dfrac{1}{2x^2}+C"


7) Substitute back "u= \dfrac{v}{x}" "\\dfrac{v}{x}\t=- \\dfrac{1}{2x^2}+C"
8) Multiply by x"v=- \dfrac{1}{2x}+Cx" 9) Substitute back "v= \dfrac{1}{y}" "\\dfrac{1}{y}\t=- \\dfrac{1}{(2x)}+Cx" "\\dfrac{1}{y}\t=- \\dfrac{1}{(2x)}+ \\dfrac{2Cx^2}{2x}= \\dfrac{(2Cx^2-1)}{2x}"

"y= \dfrac{2x}{2Cx^2-1}"

Since C is constant, for simplicity of solution we can introduce a new constantC₁ = 2C so finally"y= \dfrac{2x}{(C_1x^2-1)}"